OpenStudy (anonymous):

Where does the formula for the volume of a cylinder derive from? Give an example and provide evidence to support your claim.

OpenStudy (anonymous):

dV = rdθdrdh \[\int\limits_{0}^{r}\int\limits_{0}^{h}\int\limits_{0}^{2\pi} rdθdrdh \]

OpenStudy (xishem):

Your variables of integrations don't match up with your integrals. But otherwise, yeah, that.

OpenStudy (anonymous):

Hmm why doesn't it match? isn't theta a dummy variable?

OpenStudy (anonymous):

oh nvm i see what you're saying.

OpenStudy (anonymous):

Thank you!!!!!